# DreamDrone

### v2

@@ Abstract
We introduce DreamDrone, an innovative method for generating unbounded flythrough scenes from textual prompts. Central to our method is a novel feature-correspondence-guidance diffusion process, which utilizes the strong correspondence of intermediate features in the diffusion model. Leveraging this guidance strategy, we further propose an advanced technique for editing the intermediate latent code, enabling the generation of subsequent novel views with geometric consistency. Extensive experiments reveal that DreamDrone significantly surpasses existing methods, delivering highly authentic scene generation with exceptional visual quality. This approach marks a significant step in zero-shot perpetual view generation from textual prompts, enabling the creation of diverse scenes, including natural landscapes like oases and caves, as well as complex urban settings such as Lego-style street views. Our code is publicly available.
Abstract
In this supplementary material, we provide more comprehensive ablation studies, comparisons of visual results with text-to-video methods, and additional visual results of our method. Finally, we discuss the limitations of this approach.
Figure 1: Visualization results of DreamDrone. We visualize 30 continuous views for each prompt. The generated scenes are geometry-consistent across adjacent camera views and the details are added gradually when the camera is moving forward.
[…]
@@ 1 Introduction
Recent advances in vision and graphics have enabled the synthesis of multi-view consistent 3D scenes along extended camera trajectories, a capability previously explored in works like [17, 19, 3, 7]. This emerging task, termed perpetual view generation [19], involves synthesizing views from a flying camera along an arbitrarily long trajectory, starting from a single RGB image. Our study extends this concept, asking: Can we fly into scenes from any context, real or imagined, using textual prompts? This paper advances the novel concept of zero-shot, training-free perpetual view generation. This entails creating infinite scenes from textual prompts without optimization or fine-tuning on specific datasets. The challenge is considerable: as the camera moves forward, unseen regions must be seamlessly integrated, new details resolved as they approach the camera, all while maintaining photorealism and diversity. Existing research in video synthesis, image inpainting, and view synthesis does not fully address these requirements due to various limitations.
Recent text-to-video generation methods leverage advancements in diffusion models [2, 9, 40, 47, 14, 23] to extend their application to the temporal domain. These approaches show promise, but often require extensive training on large datasets, a costly and resource-intensive endeavor. T2V-0 [14] pioneered zero-shot text-to-video generation, producing notable results. However, limitations are evident as it can generate only a limited number of novel frames, and the quality degrades for longer videos. Additionally, these methods often overlook crucial video elements, such as scene geometry and camera movement. In contrast, text-to-3D methods like [18, 25, 30, 41, 44] utilize geometry to synthesize high-quality novel views but are constrained to limited camera motion and single-object view generation.
SceneScape [7], while conceptually similar to our work, diverges significantly in execution and capability. Its primary strategy involves outfilling regions during backward camera movement using 3D meshes. However, SceneScape encounters challenges with forward-moving cameras and in generating outdoor scenes. These issues stem from 1) difficulty in accurately filling irregular holes and preserving correspondence with original regions, 2) distortions in wrapped images remaining uncorrected, and 3) an inability to adequately detail areas enlarged by forward camera motion. Additionally, its dependence on mesh representations confines SceneScape to indoor scenarios, limiting its versatility. Our comparative analysis in Fig. 4 illustrates these limitations in SceneScape.
In this paper, we introduce DreamDrone, a novel method designed to generate infinite scenes from textual prompts. DreamDrone leverages off-the-shelf text-to-image diffusion models [34] and depth estimation models [32]. Central to our proposed method is a feature-correspondence-guidance diffusion process, specifically tailored to create geometry-consistent novel views during the denoising phase. Furthermore, we develop an innovative approach for editing the intermediate latent code, facilitating the generation of subsequent novel views. Complementing this, a cross-view self-attention module is employed to ensure consistent correspondence across adjacent views.
@@
Our experiments demonstrate that the proposed DreamDrone effectively leads to high-quality and geometry-consistent scene generation, as evidenced in our results (see Fig. 1 and Figs. 3 and 4 in Sec. 4). Notably, despite other works relying on large-scale data for training, our method exhibits comparable, even superior performance. A significant advantage of DreamDrone is its versatility: it is adept not only at generating real-world scenarios but also shows promising capabilities in creating imaginative scenes. In summary, our key contributions are as follows:
  • We propose a novel approach for zero-shot text-to-perpetual view generation, capable of producing diverse indoor and outdoor scenes in various imaginative styles. This method achieves significant results without requiring fine-tuning or optimization on specific datasets.
  • We develop a novel feature-correspondence-guidance diffusion process to maintain geometric consistency in novel view generation. This approach involves a strategic modification of the diffusion model’s intermediate features, enhancing the fidelity and detail in each newly generated view.
  • Extensive experiments demonstrate that DreamDrone can generate infinite sequences, exemplified by 30 frames in Fig. 1. It handles a broad spectrum of scenarios, from realistic to fantastical, encompassing oases, caves, and urban scenes in styles like Lego or Van Gogh.
@@ 2 Related Works
InfNat [19], InfNat-0 [17], and DiffDreamer [3] use iterative training for long-trajectory perpetual view extrapolation. InfNat [19] pioneered the perpetual view generation task with a database for infinite 2D landscapes. InfNat-0 [17] adapted this to 3D, introducing a render-refine-repeat phase for novel views. DiffDreamer [3] improved consistency with image-conditioned diffusion models. However, trained on specific datasets, these methods lack robustness in new scenarios, like urban environments, and their non-text-based conditioning restricts generative capabilities. In very recent concurrent work, SceneScape [7] generates limited scenes when the camera is moving backward, which focuses on 3D-consistent ‘zoom out’ video generation. When the camera is moving forward, there is not any specific mechanism to add details and alleviate distortion. Moreover, SceneScape can only generate limited indoor scenes because they applies a mesh as a unified representation.
@@
Several text-to-3D generation methods [1, 27, 4, 29, 49] apply text-3D pair databases to learning a mapping function. However, supervised strategies remain challenging due to the lack of large-scale aligned text-3D pairs. CLIP-based [31] 3D generation methods [51, 28, 15, 13, 12] apply pre-trained CLIP model to create 3D objects by formulating the generation as an optimization problem in the image domain. Recent text-to-3D methods like [18, 24, 25, 44, 30] blend text-to-image diffusion models [34] with neural radiance fields [26] for training-free 3D object generation. Other approaches [37, 20, 21, 33] focus on novel view synthesis from a single image, often limited to single objects or small camera motion ranges. Text2room [10] generates 3D indoor scenes from text, but is confined to room meshes. Our DreamDrone diverges by creating endless, diverse indoor and outdoor scenes, in styles ranging from Lego to Van Gogh.
@@
Generating videos from textual descriptions [23, 11, 9, 2, 36, 50, 43] poses significant challenges, primarily due to the scarcity of high-quality, large-scale text-video datasets and the inherent complexity in modeling temporal consistency and coherence. CogVideo [11] addresses this by incorporating temporal attention modules into the pre-trained text-to-image model CogView2 [6]. The video diffusion model [9] employs a space-time factorized U-Net, utilizing combined image and video data for training. Video LDM [2] adopts a latent diffusion approach for generating high-resolution videos. However, these methods typically do not account for the underlying 3D scene geometry in scene-related video generation, nor do they offer explicit control over camera movement. Additionally, their reliance on extensive training with large datasets can be prohibitively costly. While T2V-0 [14] introduced the concept of zero-shot text-to-video generation, its capability is limited to generating a small number of novel frames (up to 8 frames), with diminished quality in longer video sequences.
@@ 3 Method
Figure 2: Overview of DreamDrone. Starting from a real or generated RGBD ($I$, $D$) image at the current view, we apply DDIM backward steps to obtain intermediate latent code $x_{t_{1}}$ at timestep $t_{1}$ using a pre-trained U-Net model. A low-pass wrapping strategy is applied to generate latent code for the next novel view. A few more DDPM forward steps from timestep $t_{1}$ to $t_{2}$ are followed for enlarging the degree of freedom w.r.t. the wrapped latent code. In the reverse progress, we apply pre-trained U-Net to recover the novel view from $x_{t_{2}}^{\prime}$. The cross-view self-attention module and feature-correspondence guidance are applied to maintain consistency between $x_{t_{2}}$ and $x_{t_{2}}^{\prime}$. The right side shows the wrapped images and our generated novel view $I^{\prime}$. Our method greatly alleviates blurring, in-consistency, and distortion. The overall pipeline is zero-shot and training-free.
@@ 3.1 Preliminaries
In this paper, we implement our method based on the recent state-of-the-art text-to-image diffusion model (i.e. Stable Diffusion [34]). Stable diffusion is a latent diffusion model (LDM), which contains an autoencoder $\mathcal{D}(\mathcal{E}(\cdot)))$ and a U-Net [35] denoiser. Diffusion models are founded on two complementary random processes. The DDPM forward process, in which Gaussian noise is progressively added to the latent code of a clean image: $\boldsymbol{x}_{0}$:
$\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\boldsymbol{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{z},$ (1)
where $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\left\{\alpha_{t}\right\}$ are the noise schedule.
The backward process is aimed at gradually denoising $\boldsymbol{x_{T}}$, where at each step a cleaner image is obtained. This process is achieved by a U-Net $\boldsymbol{\epsilon}_{\theta}$ that predicts the added noise $\boldsymbol{z}$. Each step of the backward process consists of applying $\boldsymbol{\epsilon}_{\theta}$ to the current $\boldsymbol{x}_{t}$, and adding a Gaussian noise perturbation to obtain a cleaner $\boldsymbol{x}_{t-1}$.
Classifier-guided DDIM sampling [5] aims to restore images from noise conditioned on the class label. Given the diffusion model $\boldsymbol{\epsilon}_{\theta}$, the latent code $\boldsymbol{x}_{t}$ at timestep $t$, the classifier $p_{\theta}(y|x_{t})$, and the gradient scale $s$, the sampling process for obtaining $\boldsymbol{x}_{t-1}$ is formulated as:
$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\log p_{\phi}(y|\boldsymbol{x}_{t}),$ (2)
@@
$\boldsymbol{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\epsilon},$ (3)
@@
In the self-attention block of the U-Net, features are projected into queries $\mathit{\mathbf{Q}}$, keys $\mathit{\mathbf{K}}$, and values $\mathit{\mathbf{V}}$. The output of the block $\boldsymbol{o}$ is obtained by:
$\boldsymbol{o}=\mathbf{A}\mathbf{V},\;\;\textup{where}\;\mathbf{A}=\textup{Softmax}(\mathit{\mathbf{Q}}\mathit{\mathbf{K}}^{\top})$ (4)
The self-attention operation allows for long-range interactions between features.
@@ 3.2 Overview
In generating infinite scenes, the process involves two key steps. Firstly, the RGB for the forthcoming frame is calculated based on the RGBD data and camera extrinsics of the current frame. Secondly, to address the presence of black holes in the next frame’s image, inpainting [34, 22, 48] and outpainting [16, 46] techniques are employed for correction. Despite this approach, several challenges arise: 1) The irregular nature of the black holes makes it difficult for existing inpainting methods to fill these areas while maintaining semantic coherence with the rest of the image. 2) Inaccuracies in the depth data can cause distortions in certain image areas when wrapping. These distortions tend to accumulate over time with the movement of the camera, leading to a noticeable decline in image quality. 3) Additionally, as the image advances, it requires the enhancement of details in certain areas, a task at which current inpainting and outpainting methods are inadequate, particularly due to their lack of super-resolution capabilities.
To avoid generating distorted areas, ensure semantic consistency between adjacent frames, and enhance image detail during camera movement, we leverage the robustness and rich semantic capabilities of the denoise phase and decoder in the stable diffusion model. We edit the features of the diffusion based on depth and camera extrinsics, followed by denoising to produce a high-definition image for the subsequent frame. Our comprehensive design is detailed in Fig. 2. Initially, we obtain the latent code of the previous frame’s RGB image at timestep $t_{1}$ through the inverse diffusion process. We then wrap the current frame’s latent code using depth information and camera extrinsics. To maintain content and high-frequency detail consistency while wrapping the latent code, we introduce a low-pass wrapping strategy. Then several DDPM (Denoising Diffusion Probabilistic Models) processes are applied to increase the generative process’s flexibility, adding noise from timestep $t_{1}$ to $t_{2}$. Finally, we obtain the RGB image of the next frame through the denoising process. For ensuring semantic consistency between adjacent frames, we utilize a cross-frame self-attention strategy, injecting the key and value from the previous frame’s UNet network into the next frame. Furthermore, we propose a feature-correspondence-guidance strategy to influence and maintain semantic consistency between adjacent frames during the denoise latent code process. Our entire pipeline requires only a pre-trained text-to-image diffusion model and a depth estimation model, eliminating the need for any additional training or fine-tuning.
3.3 Wrapping latent codes

### v3

@@ Abstract
We introduce DreamDrone, a novel zero-shot and training-free pipeline for generating unbounded flythrough scenes from textual prompts. Different from other methods that focus on warping images frame by frame, we advocate explicitly warping the intermediate latent code of the pre-trained text-to-image diffusion model for high-quality image generation and generalization ability. To further enhance the fidelity of the generated images, we also propose a feature-correspondence-guidance diffusion process and a high-pass filtering strategy to promote geometric consistency and high-frequency detail consistency, respectively. Extensive experiments reveal that DreamDrone significantly surpasses existing methods, delivering highly authentic scene generation with exceptional visual quality, without training or fine-tuning on datasets or reconstructing 3D point clouds in advance.
Figure 1: Visualization results of DreamDrone. Given a single scene image and the textual description, our approach generates novel views corresponding to user-defined camera trajectory, without fine-tuning on any dataset or reconstructing the 3D point cloud in advance.
@@ 1 Introduction
Recent advances in vision and graphics have enabled the synthesis of multi-view consistent 3D scenes along extended camera trajectories [18, 20, 3, 7]. This emerging task, termed perpetual view generation [20], involves synthesizing views from a flying camera along an arbitrarily long trajectory, starting from a single RGBD image.
Previous methodologies predominantly engage in warping images frame by frame with traditional 3D geometric knowledge when given RGBD images and subsequent camera extrinsic. However, this operation often leads to blurriness and distortion in images, which arises from inaccurate interpolation, the mismatch between discrete pixels and continuous transformations, and inaccurate depth data. Moreover, such blurriness and distortion tend to amplify with the accumulation of warp operations.
To further alleviate the errors caused by frame-by-frame warp operations, two primary paths have been proposed. i) Some methods [18, 20, 3] try to train a refiner on natural scene datasets. The advantage of this frame-by-frame approach is that it allows for arbitrary changes in camera trajectory during the scene generation process, offering users a higher degree of freedom and enabling infinite generation. However, this training-based method can only be used in natural scenes and cannot be generalized to arbitrary indoor/outdoor scenes or scenes of various styles. ii) Another solution is to first reconstruct the 3D scene model using text prompts, then render 2D RGB images according to the camera trajectory [7, 52, 10]. Although this solution yields more coherent 2D image sequences, the quality of the rendered images highly depends on the quality of the 3D scene model. This method cannot guarantee good rendering effects from every viewpoint. Additionally, since this method requires the reconstruction of 3D point clouds, it cannot achieve "infinite" scene generation in the same way as the frame-by-frame strategy.
In this paper, we advocate that a more general and flexible perpetual view generation pipeline should possess the following capabilities:
i) are versatile across diverse scenes, including indoor and outdoor scenes, as well as scenes depicted in various styles; ii) allow users to interactively control the camera trajectory during the process of scene generation, while ensuring the high quality of the generated images and the semantic consistency between adjacent frames; and iii) enable seamless transitions from one scene to another.
To this end, we introduce DreamDrone, a novel zero-shot, training-free, infinite scene generation pipeline from text prompts, which does not require any optimization or fine-tuning on any dataset. A core principle of our approach is to warp the latent code of a pre-trained text-to-image diffusion model rather than the frames, enriching it with temporal and geometric consistency. To be specific, given RGBD image $I$ of the current view and camera rotation $\mathcal{R}$ and translation $\mathcal{T}$ for the next view (which is interactively defined by users), we first obtain the latent code $x_{t_{1}}$ of the diffusion model at timestep $t_{1}$, warp it to latent code of the next view $x^{\prime}_{t_{1}}$ based on $\mathcal{R}$ and $\mathcal{T}$, and denoise $x^{\prime}_{t_{1}}$ to the image $I^{\prime}$ of the next view. To ensure geometry consistency across adjacent views, we propose a novel feature-correspondence-guidance diffusion process when denoising from $x^{\prime}_{t_{1}}$ to image at the next view $I^{\prime}$. Moreover, we propose a novel high-pass filter mechanism when warping the latent code $x_{t_{1}}$, for preserving high-frequency details across adjacent views.
@@
Our experiments demonstrate that the proposed DreamDrone effectively leads to high-quality and geometry-consistent scene generation. Quantitative and qualitative results demonstrate our comparable, even superior performance compared with other training-based and training-free methods from the aspects of temporal consistency and image quality. Moreover, the significant advantage of DreamDrone is its versatility: it is adept not only at generating real-world scenarios but also shows promising capabilities in creating imaginative scenes. Additionally, users can interactively control the camera trajectory (Fig. 5) and shuttle from one scene to another (Fig. 4). Our contributions are summarized as follows:
  • To our best knowledge, we are the first attempt to generate novel views by explicitly warping the latent code of the pre-trained diffusion model.
  • A novel feature-correspondence-guidance diffusion process is proposed to enforce geometry consistency across adjacent views. Moreover, a high-pass filtering strategy is introduced to preserve high-frequency details for novel views.
  • Extensive experiments demonstrate that our method generates high-quality and geometry-consistent novel views for any scene, from realistic to fantastical. More interestingly, our method realizes the scene shuttle, i.e., travels from one scene to another when the user controls the camera trajectory.
@@ 2 Related Works
Perpetual view generation extrapolates unseen content outside a single image. InfNat [20], InfNat-0 [18], and DiffDreamer [3] use iterative training for long-trajectory perpetual view extrapolation. InfNat [20] pioneered the perpetual view generation task with a database for infinite 2D landscapes. InfNat-0 [18] adapted this to 3D, introducing a render-refine-repeat phase for novel views. DiffDreamer [3] improved consistency with image-conditioned diffusion models. However, these methods lack robustness in complex and urban environments. In very recent concurrent work, SceneScape [7] and WonderJourney [52] firstly generate 3D point cloud for scene by zoom-out and inpainting strategy. 2D image sequences are further rendered based on the reconstructed 3D point cloud. However, the accuracy of the 3D model critically impacts performance, particularly with novel camera trajectories.
@@
Several text-to-3D generation methods [1, 28, 4, 30, 54, 15] apply text-3D pair databases to learning a mapping function. However, supervised strategies remain challenging due to the lack of large-scale aligned text-3D pairs. CLIP-based [33] 3D generation methods [56, 29, 16, 13, 12] apply pre-trained CLIP model to create 3D objects by formulating the generation as an optimization problem in the image domain. Recent text-to-3D methods like [19, 25, 26, 47, 31, 38, 50] blend text-to-image diffusion models [36] with neural radiance fields [27] for training-free 3D object generation. Other approaches [41, 21, 22, 35] focus on novel view synthesis from a single image, often limited to single objects or small camera motion ranges. Text2room [10] generates 3D indoor scenes from text prompts, but is confined to room meshes.
@@
Generating videos from textual descriptions [24, 11, 9, 2, 39, 55, 46, 40] poses significant challenges, primarily due to the scarcity of high-quality, large-scale text-video datasets and the inherent complexity in modeling temporal consistency and coherence. CogVideo [11] addresses this by incorporating temporal attention modules into the pre-trained text-to-image model CogView2 [6]. The video diffusion model [9] employs a space-time factorized U-Net, utilizing combined image and video data for training. Video LDM [2] adopts a latent diffusion approach for generating high-resolution videos. However, these methods typically do not account for the underlying 3D scene geometry in scene-related video generation, nor do they offer explicit control over camera movement. Additionally, their reliance on extensive training with large datasets can be prohibitively costly. While T2V-0 [14] introduced the concept of zero-shot text-to-video generation, its capability is limited to generating a small number of novel frames, with diminished quality in longer video sequences.
@@ 3 Method
We formulate the task of perpetual view generation as follows: given a starting image $I$, we generate the next view image $I^{\prime}$ corresponding to an arbitrary camera pose $\left\{\mathcal{R},\mathcal{T}\right\}$, where the camera pose can be specified or via user’s control.
@@ 3.1 Preliminaries
We implement our method based on the recent state-of-the-art text-to-image diffusion model (i.e. Stable Diffusion [36]). Stable diffusion is a latent diffusion model (LDM), which contains an autoencoder $\mathcal{D}(\mathcal{E}(\cdot)))$ and a U-Net [37] denoiser. Diffusion models are founded on two complementary random processes. The DDPM forward process, in which Gaussian noise is progressively added to the latent code of a clean image: $\bm{x}_{0}$:
$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sqrt{1-\alpha_{t}}\bm{z},$ (1)
where $\bm{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\left\{\alpha_{t}\right\}$ are the noise schedule.
The backward process is aimed at gradually denoising $\bm{x_{T}}$, where at each step a cleaner image is obtained. This process is achieved by a U-Net $\bm{\epsilon}_{\theta}$ that predicts the added noise $\bm{z}$. Each step of the backward process consists of applying $\bm{\epsilon}_{\theta}$ to the current $\bm{x}_{t}$, and adding a Gaussian noise perturbation to obtain a cleaner $\bm{x}_{t-1}$.
Classifier-guided DDIM sampling [5] aims to generate images from noise conditioned on the class label. Given the diffusion model $\bm{\epsilon}_{\theta}$, the latent code $\bm{x}_{t}$ at timestep $t$, the classifier $p_{\theta}(y|x_{t})$, and the gradient scale $s$, the sampling process for obtaining $\bm{x}_{t-1}$ is formulated as:
$\hat{\epsilon}=\bm{\epsilon}_{\theta}(\bm{x}_{t})-\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\bm{x}_{t}}\log p_{\phi}(y|\bm{x}_{t}),$ (2)
@@
$\bm{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\frac{\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}}\hat{\epsilon},$ (3)
@@
In the self-attention block of the U-Net, features are projected into queries $\mathit{\mathbf{Q}}$, keys $\mathit{\mathbf{K}}$, and values $\mathit{\mathbf{V}}$. The output of the block $\bm{o}$ is obtained by:
$\bm{o}=\mathbf{A}\mathbf{V},\;\;\textup{where}\;\mathbf{A}=\textup{Softmax}(\mathit{\mathbf{Q}}\mathit{\mathbf{K}}^{\top})$ (4)
The self-attention operation allows for long-range interactions between image tokens.
@@ 3.2 Overview
Figure 2: Overview of our proposed pipeline. Starting from a real or generated RGBD ($I$, $D$) image at the current view, we apply DDIM inversion to obtain intermediate latent code $x_{t_{1}}$ at timestep $t_{1}$ using a pre-trained U-Net model. A warping with the high-pass filter strategy is applied to generate latent code for the next novel view. A few more DDPM forward steps from timestep $t_{1}$ to $t_{2}$ are applied for enlarging the degree of freedom w.r.t. the warped latent code. In the denoising process, we apply pre-trained U-Net to generate the novel view from $x_{t_{2}}^{\prime}$. The cross-view self-attention module and feature-correspondence guidance are applied to maintain the geometry correspondence between $x_{t_{2}}$ and $x_{t_{2}}^{\prime}$. The right side shows the warped image and our generated novel view $I^{\prime}$. Our method greatly alleviates blurring, inconsistency, and distortion. The overall pipeline is zero-shot and training-free.
Perpetual view generation as the camera moves presents a complex challenge. This process involves seamlessly filling in unseen regions caused by image warping, adding details to objects as they come closer, while ensuring the imagery remains realistic and diverse. Prior works [18, 20, 3] have focused on training a refiner to enhance details and create new content for areas requiring inpainting or outpainting. These efforts have shown promising outcomes, yet the effectiveness of the refiner is generally limited to scenarios that align with the training dataset.
Since diffusion models can generate high-quality large-variety images from random latent code, a direct solution arises: can we modify the powerful pre-trained text-to-image diffusion model as a refiner? Empirically, DDIM inversion strategy [14, 43] can obtain the intermediate latent code at each timestep and the image can be reconstructed by those latent codes. To this end, we attempt to explicitly warp the latent code of the current view and generate the novel view by the pre-trained text-to-image diffusion model.
Our overall pipeline is illustrated in Fig. 2. Initially, we obtain the latent code $x_{t_{1}}$ of the current view’s RGB image $I$ at timestep $t_{1}$ through the DDIM inversion process. We then warp the current frame’s latent code $x_{t_{1}}$ to the next view $x^{\prime}_{t_{1}}$ using depth information and camera extrinsic parameters. However, directly denoising from $x^{\prime}_{t_{1}}$ to the image also suffers from blurry, which results in the non-integer pixel coordinates and the interpolation operation. More noise is added from $x^{\prime}_{t_{1}}$ at timestep $t_{1}$ to $x^{\prime}_{t_{2}}$ at timestep $t_{2}$ by DDPM forward operation, for generating high-quality images. The side effect of DDPM is the geometry inconsistency between adjacent views. To this end, we propose a feature-correspondence-guidance denoising strategy to enforce geometry consistency. Moreover, a high-pass filtering strategy is proposed to maintain the consistency of high-frequency details between adjacent views. Please refer to Fig. 3 for the motivation of our proposed modules. Our overall pipeline requires only a pre-trained text-to-image diffusion model and a depth estimation model, eliminating the need for any additional training or fine-tuning.
3.3 Warping latent codes